OpenStudy (anonymous):

an accepted relationship between stopping distance, d in feet, and the speed of a car, in mph, is d(v)=1.2v+o.o4v^2 on dry, level concrete. a) how many feet will it take a car traveling 50 mph to stop on dry, level concrete? b) if an accident occurs 250 feet ahead, what is the maximum speed at which one can travel to avoid being involved in the accident?

OpenStudy (blurbendy):

so what happens if you put 50 in for v in the equation

OpenStudy (anonymous):

it would be 160

OpenStudy (anonymous):

how would you do b??

OpenStudy (whpalmer4):

\[d(v) = 250 = 1.2v+0.04v^2\]Can you solve that for\(v\)?

OpenStudy (anonymous):

are we solving for v?

OpenStudy (whpalmer4):

You can use the quadratic formula to find the solutions of a polynomial\(ax^2+bx+c=0, a\ne0\): \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] One of the solutions won't make sense here.

OpenStudy (whpalmer4):

Yes, solve for \(v\).

OpenStudy (anonymous):

would the answer be -.04 +- square root of 1200.0016)/2.4

OpenStudy (whpalmer4):

What are your values of a, b, and c?

OpenStudy (anonymous):

opps i mean -1.2 +- squareroot of (41.44)/.08

OpenStudy (whpalmer4):

What part of that are you dividing by 0.08?

OpenStudy (anonymous):

all of them

OpenStudy (whpalmer4):

That's what I figured, but that's not what you wrote. What you wrote is \[-1.2\pm \frac{\sqrt{41.44}}{0.08}\]

OpenStudy (whpalmer4):

What are your numeric answer?

OpenStudy (whpalmer4):

answers

OpenStudy (whpalmer4):

\[(-1.2 + \sqrt{41.44})/0.08 =\] \[(-1.2-\sqrt{41.44})/0.08 = \]

OpenStudy (whpalmer4):

The negative value of \(v\) doesn't make sense as an answer in this context, so we'll discard it.

OpenStudy (anonymous):

first would be -95.47 second would be 65.47 would the answer be the second one since it is the max.

OpenStudy (whpalmer4):

Yes, the second one is the only one that makes sense, and that's the correct answer.